# Harness as an Asset: Enforcing Determinism via the Convergent AI Agent Framework (CAAF)

### 18 Apr 2026 v1

@@ Abstract
Large Language Models (LLMs) produce a controllability gap in safety-critical engineering: even low rates of undetected constraint violations render a system undeployable. Current orchestration paradigms suffer from sycophantic compliance, context attention decay (Liu et al., 2024), and stochastic oscillation during self-correction (Huang et al., 2024).
We introduce the Convergent AI Agent Framework (CAAF), which transitions agentic workflows from open-loop generation to closed-loop Fail-Safe Determinism via three pillars: (1) Recursive Atomic Decomposition with physical context firewalls; (2) Harness as an Asset, formalizing domain invariants into machine-readable registries enforced by a deterministic Unified Assertion Interface (UAI); and (3) Structured Semantic Gradients with State Locking for monotonic convergence.
Empirical evaluation across two domains—SAE Level 3 (L3) autonomous driving (AD) (n=30n{=}30, 7 conditions) and pharmaceutical continuous flow reactor design (n=20n{=}20, 4 conditions including a Mono+UAI ablation)—shows that CAAF-all-GPT-4o-mini achieves 100% paradox detection while monolithic GPT-4o achieves 0% (even at temperature==0). The pharmaceutical benchmark features 7 simultaneous constraints with nonlinear Arrhenius interactions and a 3-way minimal unsatisfiable subset, representing a structurally harder challenge than the 2-constraint AD paradox. Alternative multi-agent architectures (debate, sequential checking) also achieve 0% across 80 trials, confirming that CAAF’s reliability derives from its deterministic UAI, not from multi-agent orchestration per se. A Mono+UAI ablation (95%) isolates UAI as the core contribution. CAAF’s reliability is invariant to prompt hints; all components use a single commodity model, enabling fully offline deployment.
11footnotetext: This work was conducted in the author’s personal capacity as an independent researcher. The views expressed are those of the author alone and do not represent any current or former employer. All domain examples are constructed from publicly available physics and engineering literature; no proprietary or confidential information was used. CAAF is a general-purpose framework not tied to any specific organization’s products or services.
@@ 1 Introduction
While techniques such as Chain-of-Thought (Yao et al., 2023) and iterative self-reflection have marginally improved LLM reliability, they remain vulnerable to two well-documented structural failure modes: Context Rot, where models deprioritize safety constraints introduced early in long prompts (Liu et al., 2024); and Stochastic Oscillation, where models oscillate between constraint violations when simultaneously satisfying competing requirements (Huang et al., 2024).
This paper introduces the Convergent AI Agent Framework (CAAF), which integrates emerging industrial practices of Harness Engineering (Böckeler, 2024) with classical Cybernetic Control Theory to address these failure modes. CAAF builds on insights from three research directions: (a) evidence that LLMs cannot reliably self-correct intrinsic reasoning failures without external grounding (Huang et al., 2024); (b) the principle that domain invariants can be formalized as executable machine contracts (Meyer, 1992; CNCF, 2019); and (c) the observation that textual feedback signals can be structured as analogue gradients to guide iterative improvement (Yuksekgonul et al., 2024).
@@
This paper validates CAAF through two independent domain benchmarks—automotive engineering and pharmaceutical continuous flow reactor design—and demonstrates that alternative multi-agent architectures (debate, sequential checking) fail to achieve the same reliability on these tasks. We argue, based on the architectural decomposition rather than on domain count alone, that CAAF addresses a general structural problem: any workflow in which an LLM generates a complex artifact that must simultaneously satisfy multiple formally verifiable constraints. The domain-specific component is exclusively the Harness Registry and its UAI validators; the decomposition engine, convergence loop, and paradox detection mechanism are designed to be domain-agnostic. We therefore conjecture that wherever constraints can be expressed as executable assertions—whether physical laws, regulatory requirements, financial models, or infrastructure Service-Level Agreements (SLAs)—the same architectural pattern applies; broader empirical validation beyond the two reported domains remains future work (Section 11.6). In this sense, we propose CAAF as a candidate constraint satisfaction layer for LLM-generated artifacts: analogous to how a type system catches errors at compile time before code reaches production, CAAF is designed to catch constraint violations at generation time before engineering artifacts reach downstream processes.
@@ 3.2 Pillar 2: Harness as an Asset (HaaA)
3.3 Pillar 3: Structured Semantic Gradients and Monotonic Convergence
@@
Monotonic Convergence via State Locking: To prevent stochastic oscillation (“fixing A while breaking B”), CAAF implements State Locking. When a dimension is validated as PASS, its corresponding JSON schema node is marked read_only: true in the subsequent iteration prompt. This forces the Executor to concentrate creative inference only on failing dimensions, guaranteeing that the set of verified constraints grows monotonically across iterations (Eq. 2 below). Importantly, State Locking operates within each Executor’s scope: because nodes are context-firewalled, locking a verified dimension in one node does not constrain the adjustment of independent dimensions in sibling nodes. The Context Firewall and State Locking are complementary mechanisms—the former prevents cross-node contamination, while the latter prevents intra-node regression.
Formal Convergence Property. Let $\mathcal{C}=\{c_{1},\ldots,c_{K}\}$ denote the set of modeled constraints, and let $V_{t}\subseteq\mathcal{C}$ denote the set of constraints satisfied (verified) at iteration $t$. Under CAAF’s State Locking mechanism:
@@
That is, the set of verified constraints grows monotonically. This property holds under two assumptions: (a) the Harness is internally consistent (no contradictory assertions), and (b) UAI evaluations are deterministic. The loop terminates in one of two states: $V_{T}=\mathcal{C}$ (all constraints satisfied, PASS) or the Reviewer determines that $\mathcal{C}$ is unsatisfiable (FAILED_PARADOX). The monotonicity guarantee follows directly from State Locking: once $c_{i}\in V_{t}$, its corresponding schema field is marked read_only, preventing the Executor from modifying it in subsequent iterations.
@@
Figure 5: Stochastic oscillation (naive reflection, n=20n{=}20) vs. deterministic termination (CAAF). (a) Each dot is one run’s proposed speed at a given iteration; the naive baseline oscillates predominantly between the two constraint boundaries (v=55v{=}55 and $84$ km/h), with occasional outlier speeds. (b) The same data viewed as constraint-violation count: all naive runs remain below the axis (paradox undetected, predominantly 1 violation per iteration with occasional 2), while CAAF halts at iteration 1 with both constraints identified as irreconcilable. No naive run converges; CAAF requires zero iterative search.
@@ 4.2 Experimental Design and Results
Table 1: Physical Paradox Detection — Full Experimental Results (n=30n{=}30)
@@ 4.5 PASS-Path Convergence on the Satisfiable Variant
Monotone-convergence runs (Eq. 2) 20/20
@@ 5.2 Monolithic Failure: Systematic Safety-Priority Collapse
System Test Condition Paradox Detected Failure Distribution Mono GPT-4o (t=0) Clean (safety-only) 0/20 Rear-Viol: 18, Fwd-Viol: 2 Mono GPT-4o (t=0) Noise (VIP+thermal) 0/20 Rear-Viol: 14, Fwd-Viol: 6 CAAF-all-mini Clean (safety-only) 20/20 CAAF-all-mini Noise (VIP+thermal) 20/20
@@ 7.2 Experimental Results
We evaluated four conditions (n=20n{=}20 each) using GPT-4o-mini, adding a Mono+UAI (ReAct (Yao et al., 2023)) ablation condition to isolate the contribution of the deterministic UAI tool:
Table 7: Pharmaceutical Flow Reactor Paradox Results (n=20n{=}20 per condition)
# System Hint $n$ Correct% Failure Mode Distribution 1 Monolithic GPT-4o-mini × 20 0% Impurity-Viol: 10, Conv-Viol: 7, JSON: 3 2 Monolithic GPT-4o-mini 20 65% Impurity-Viol: 6, JSON: 1 3 CAAF-all-mini × 20 100% Paradox detected: 20/20 4 Mono + UAI (ReAct) × 20 95% Impurity-Viol: 1
@@
Figure 11: Failure mode distribution across six experimental conditions (n=20n{=}20 each) in the pharmaceutical domain. Baselines exhibit diverse failure modes dominated by impurity and conversion violations, with the sequential baseline additionally exhibiting a high JSON-failure rate (6/20) after multiple retry cycles. CAAF achieves perfect paradox detection; Mono+UAI’s single failure (1/20) is an impurity constraint omission that RAD’s atomic decomposition is designed to prevent (not directly ablated; see Finding 8).
@@ 7.3 Analysis
Finding 7 — Hint Fragility Scales with Complexity. In the AD domain (2-constraint paradox), the hint condition achieved 100% detection. In the pharma domain (3-way paradox with nonlinear interactions), hint-aided detection drops to 65%. This confirms that prompt engineering reliability degrades as constraint complexity increases—precisely the regime where architectural enforcement becomes most critical.
Finding 8 — Mono+UAI Ablation: UAI as the Core Contribution. The Mono+UAI condition strips away RAD and Structured Semantic Gradients, giving a single monolithic LLM direct access to the UAI tool via ReAct-style function calling. At 95% (19/20), this ablation demonstrates that the deterministic UAI assertion engine provides the majority of CAAF’s value. The single failure (1/20) was an impurity constraint violation: the model selected $T{=}130\,^{\circ}$C, $\tau{=}60$ s but failed to detect that the resulting impurity fraction exceeded the 2% limit—an omission that RAD’s atomic decomposition is designed to prevent by assigning impurity verification to a dedicated node, though we did not ablate this specific combination (Mono+UAI+RAD).
@@
  • UAI (Harness as an Asset) provides the deterministic constraint boundary—the core reliability mechanism.
@@
  • Structured Semantic Gradients with State Locking guarantee monotonic convergence across iterations, preventing oscillation between constraint fixes.
In the current 7-constraint benchmark, a capable model with UAI access can reason through the constraint interactions in a single pass. We hypothesize that RAD’s contribution will become critical at higher constraint counts (15+) and longer specifications, where context window limitations degrade single-pass reasoning. This represents a key axis for future experimental validation.

### 25 Apr 2026 v2

@@ Abstract
Large Language Models produce a controllability gap in safety-critical engineering: even low rates of undetected constraint violations render a system undeployable. Current orchestration paradigms suffer from sycophantic compliance, context attention decay (Liu et al., 2024), and stochastic oscillation during self-correction (Huang et al., 2024).
We introduce the Convergent AI Agent Framework (CAAF), which transitions agentic workflows from open-loop generation to closed-loop fail-safe determinism via three pillars: (1) Recursive Atomic Decomposition (RAD) with physical context firewalls; (2) Harness as an Asset, formalizing domain invariants into machine-readable registries enforced by a deterministic Unified Assertion Interface (UAI); and (3) Structured Semantic Gradients with State Locking for monotonic non-regression.
This paper makes two core claims. First, an industrialization thesis: once domain invariants are formalized as an executable Harness, the Harness itself becomes a first-class enterprise asset that compounds in value as foundation models commoditize, and CAAF’s ability to deliver its reliability on commodity-tier models makes fully self-hosted, on-premises deployment architecturally feasible for regulated sectors where cloud APIs are not an option. Second, an architectural claim supported by ablation: CAAF’s three pillars address complementary failure surfaces—context isolation, deterministic grounding, and oscillation control—and none alone closes the controllability gap at commodity cost. The paper proposes no new training recipe, architecture, or scaling study; the contribution is entirely at the orchestration and industrialization layer. Evidence across two structurally complementary benchmarks (autonomous driving and pharmaceutical reactor design), three-tier UAI ablations, multi-agent baselines, and a closed-source commodity family replicated by two independent open-weight families, is reported in the body.
@@ 1 Introduction
While techniques such as Chain-of-Thought (Wei et al., 2022), ReAct (Yao et al., 2023), and iterative self-reflection have marginally improved LLM reliability, they remain vulnerable to two well-documented structural failure modes: Context Rot, where models deprioritize safety constraints introduced early in long prompts (Liu et al., 2024); and Stochastic Oscillation, where models oscillate between constraint violations when simultaneously satisfying competing requirements (Huang et al., 2024).
This paper introduces the Convergent AI Agent Framework (CAAF), which integrates emerging industrial practices of Harness Engineering (Böckeler, 2026) with classical Cybernetic Control Theory to address these failure modes. CAAF builds on insights from three research directions: (a) evidence that LLMs cannot reliably self-correct intrinsic reasoning failures without external grounding (Huang et al., 2024); (b) the principle that domain invariants can be formalized as executable machine contracts (Meyer, 1992; CNCF, 2019); and (c) the observation that textual feedback signals can be structured as analogue gradients to guide iterative improvement (Yuksekgonul et al., 2024).
@@
This paper validates CAAF through two independent domain benchmarks—automotive engineering and pharmaceutical continuous flow reactor design—and demonstrates that alternative multi-agent architectures (debate, sequential checking) fail to achieve the same reliability on these tasks. We argue, based on the architectural decomposition rather than on domain count alone, that CAAF addresses a general structural problem: any workflow in which an LLM generates a complex artifact that must simultaneously satisfy multiple formally verifiable constraints. The domain-specific component is exclusively the Harness Registry and its UAI validators; the decomposition engine, convergence loop, and paradox detection mechanism are designed to be domain-agnostic. We therefore conjecture that wherever constraints can be expressed as executable assertions—whether physical laws, regulatory requirements, or financial models—the same architectural pattern applies; broader empirical validation beyond the two reported domains remains future work (Section 11.6). In this sense, we propose CAAF as a candidate constraint satisfaction layer for LLM-generated artifacts: analogous to how a type system catches errors at compile time before code reaches production, CAAF is designed to catch constraint violations at generation time before engineering artifacts reach downstream processes.
Contributions.
This paper makes two core claims; the remaining sections supply the supporting evidence.
  1. 1.
    An industrialization thesis: the Harness as a first-class enterprise asset, and the self-hosted commodity-model deployment it unlocks. The central claim is not about model intelligence—we propose no training recipe, no new architecture, no scaling study. It is about where domain knowledge should live in an AI-augmented engineering organization. Pulling domain invariants out of prompts and into a frozen YAML+UAI contract, governed by a Change Approval Board, turns the Harness Registry into a versioned, auditable corporate asset whose value compounds as the inference layer underneath it commoditizes. Because CAAF’s reliability does not depend on frontier-model capability, all components can run on a single commodity-tier model on commodity hardware inside the enterprise firewall—which in regulated sectors (ISO 26262, ISO 13485, IEC 62443, SOTIF) is the difference between being unable to deploy AI-augmented requirements engineering at all and deploying it with the controllability contract kept in-house.
  2. 2.
    An architectural claim, validated by ablation: CAAF’s three pillars address complementary failure surfaces, and none alone closes the controllability gap at commodity cost. RAD + Context Firewall defeats Context Rot; UAI + Harness supplies the deterministic constraint boundary; Structured Semantic Gradients + State Locking prevent stochastic oscillation. A three-tier UAI ablation makes the interaction concrete: deterministic UAI grounding is necessary—monolithic and multi-agent baselines without it fail across all domains tested—but not sufficient at commodity reasoning cost, where faithful tool-call UAI access still collapses into oscillation until State Locking is added. Frontier reasoning can substitute for Pillar 3, but only at an order-of-magnitude higher per-trial cost.
The evidence supporting these two claims—two structurally complementary benchmarks (a 2-constraint L3 AD paradox and a 7-constraint pharmaceutical flow-reactor paradox with a 3-way minimal unsatisfiable subset), three-tier UAI ablations, debate / sequential-checker baselines, cross-family open-weight replication, and the full per-trial cost analysis—is reported in Sections 4–9.
@@ 2.2 Context Rot and Attention Decay
Frontier evidence (2026). We empirically confirm in Section 5.3 that context rot persists at the 2026 reasoning frontier: under cross-domain noise injection, Anthropic Claude Opus 4 with adaptive high-effort thinking (n=20n{=}20) commits the same forward-safety violation in 20/20 trials with zero variance in chosen speed, despite mean encrypted-thinking signature lengths exceeding 6,000 characters. Increased deliberation capacity does not cure context-rot failures; the reliability gap remains architectural.
@@ 3.2 Pillar 2: Harness as an Asset (HaaA)
3.3 Pillar 3: Structured Semantic Gradients and Monotonic Non-Regression
@@
Monotonic Non-Regression via State Locking: To prevent stochastic oscillation (“fixing A while breaking B”), CAAF implements State Locking. When a dimension is validated as PASS, its corresponding JSON schema node is marked read_only: true in the subsequent iteration prompt. This forces the Executor to concentrate creative inference only on failing dimensions, guaranteeing that the set of verified constraints grows monotonically across iterations (Eq. 2 below). Importantly, State Locking operates within each Executor’s scope: because nodes are context-firewalled, locking a verified dimension in one node does not constrain the adjustment of independent dimensions in sibling nodes. The Context Firewall and State Locking are complementary mechanisms—the former prevents cross-node contamination, while the latter prevents intra-node regression.
Monotonic Non-Regression + Finite Termination. Let $\mathcal{C}=\{c_{1},\ldots,c_{K}\}$ denote the set of modeled constraints, and let $V_{t}\subseteq\mathcal{C}$ denote the set of constraints satisfied (verified) at iteration $t$. Under CAAF’s State Locking mechanism:
@@
That is, the set of verified constraints is monotonically non-decreasing—once a constraint is verified, State Locking prevents the Executor from regressing it in subsequent iterations. This property holds under two assumptions: (a) the Harness is internally consistent (no contradictory assertions), and (b) UAI evaluations are deterministic. The loop terminates in one of two states: $V_{T}=\mathcal{C}$ (all constraints satisfied, PASS) or the Reviewer determines that $\mathcal{C}$ is unsatisfiable (FAILED_PARADOX). Termination is enforced by a bounded retry budget (max_iters) rather than by a formal fixed-point theorem; Eq. 2 is a non-regression invariant, not a finite-time convergence proof. We do not claim a formal contraction bound on how fast $V_{t}$ grows toward $\mathcal{C}$ in the general case; the empirical convergence behavior is reported in Sections 4.5 and 7.
@@ 4.2 Experimental Design and Results
Methodology: parse-failure exclusion.
Across all experimental cells reported in this paper, we apply a uniform parse-retry policy: any trial whose model output cannot be parsed into the required JSON schema is re-attempted with a fresh API call (max 3 attempts). Trials whose three attempts all fail are excluded from the valid-$n$ denominator and reported separately. The justification is operational: a parse-failed output is trivially detected by any production system (the JSON parser raises immediately) and cannot propagate as an undetected violation, so it does not contribute to the controllability gap this paper measures. Conflating parse failures with compliant hallucinations (well-formed JSON that violates physics) is a category error. Per-cell exclusion counts are reported wherever non-zero.
@@
Table 1: Physical Paradox Detection — Full Experimental Results (n=30n{=}30). For monolithic rows, Temp is the single-model sampling temperature; for CAAF rows, “0.7/0.0” denotes Executor/Reviewer temperatures respectively.
@@ 4.5 PASS-Path Convergence on the Satisfiable Variant
Monotone non-regression runs (Eq. 2) 20/20
@@ 5.2 Monolithic Failure: Systematic Safety-Priority Collapse
System Test Condition Paradox Detected Failure Distribution Mono GPT-4o (t=0) Clean (safety-only) 0/20 Rear-Viol: 18, Fwd-Viol: 2 Mono GPT-4o (t=0) Noise (VIP+thermal) 0/20 Rear-Viol: 14, Fwd-Viol: 6 Mono Opus 4 thinking (effort=high) Noise (VIP+thermal) 0/20 Rear-Viol: 20 (zero-variance v=55v{=}55; 11/20 also jerk-viol) CAAF-all-mini Clean (safety-only) 20/20 CAAF-all-mini Noise (VIP+thermal) 20/20
@@ 5.3 CAAF Defense: Context Firewall via RAD
5.4 Under Frontier Reasoning (CR1)
To test whether increased deliberation capacity alleviates context rot, we replicated the noisy condition using Anthropic Claude Opus 4 with adaptive high-effort extended thinking (n=20n{=}20, effort=high). Detection remained 0/20. Every trial committed $v_{t5}{=}55$ km/h (zero-variance lock onto the forward-safety boundary), violating rear safety. The mean encrypted-thinking signature length was 6,269 characters (max 13,444), confirming substantial deliberation capacity was deployed without effect on the failure outcome. Notably, 11/20 trials also violated the injected jerk constraint (>1.5>1.5 m/s3), indicating that the thinking budget was partially absorbed by the irrelevant “VIP Sleep Mode” requirement without producing a correct global trade-off.
Compared with GPT-4o (whose noisy-condition speed distribution shifted from 57.9 to 63.7 km/h mean), Opus 4 thinking is more attractor-locked under noise rather than less. This refutes the hypothesis that increased deliberation capacity cures context-rot failures: frontier reasoning with no external deterministic grounding concentrates onto the dominant local attractor more rigidly, not less. The reliability gap is architectural, not a function of raw reasoning budget. Raw data: smoke_v2_20260423_084122.
@@ 6.1 The Seesaw Effect of Naive Reflection
Figure 9: Stochastic oscillation (naive reflection, n=20n{=}20) vs. deterministic termination (CAAF). Each orange star is one naive-reflection run’s proposed speed at a given iteration; the green $\times$ marks the CAAF halt state. The naive baseline oscillates between the two constraint boundaries (v=55v{=}55 and $84$ km/h) with occasional outliers, and never declares a paradox (callout: each iteration still violates $\geq 1$ harness constraint across all 100 iteration-samples). CAAF instead exits on iteration 1 of every run by flagging both constraints as simultaneously irreconcilable and proposing no speed (20/20; callout). No naive run converges; CAAF requires zero iterative search.
@@ 7.2 Experimental Results
We evaluated eight conditions (n=20n{=}20 each) spanning two executor families (OpenAI GPT-4o-mini and Anthropic Claude). Conditions 1–4 test GPT-4o-mini with and without hints, the CAAF pipeline, and a Mono+UAI (prompt-sim) ablation in which the model is told to mentally apply the harness formulas and retry. Conditions 5–8 add 2026-frontier reasoning baselines and a cross-vendor CAAF cell: monolithic Opus 4 with adaptive high-effort thinking (Cond. 5), the same model granted true tool-call UAI access via the Anthropic tool-use API (Cond. 6), commodity-tier Haiku 4.5 with the same true tool-call UAI interface but no extended thinking (Cond. 7), and the full CAAF pipeline with all-Haiku-4.5 executors (Cond. 8).
Table 7: Pharmaceutical Flow Reactor Paradox Results (n=20n{=}20 per condition). Conditions 1–4 use GPT-4o-mini; conditions 5–8 use Anthropic Claude. The uniform parse-retry policy (§4.2) applies to every cell.
# System Hint $n$ Correct% Failure Mode Distribution 1 Monolithic GPT-4o-mini × 20 0% Impurity-Viol: 11, Conv-Viol: 9 2 Monolithic GPT-4o-mini 20 70% Impurity-Viol: 6 3 CAAF-all-mini × 20 100% Paradox detected: 20/20 4 Mono ++ UAI (prompt-sim) × 20 95% Impurity-Viol: 1 (at $T{=}130\,^{\circ}$C, $\tau{=}60$ s) 5 Mono (reasoning) Opus 4 thinking, effort=high × 20 0% Impurity: 12, Conv: 2, Dual: 6 (dual attractor at $\tau{\approx}157$ s and $\tau{=}120$ s) 6 Mono ++ UAI (true tool-call) Opus 4 thinking × 20 100% MUS $\{C1,C2,C4\}$: 20/20 (mean 3.15 iters) 7 Mono ++ UAI (true tool-call) Haiku 4.5, no thinking × 20 0% Oscillation, max_iters 20/20 (Imp: 14, Conv: 12 overlap) 8 CAAF-all-Haiku-4.5 × 20 100% Paradox detected: 20/20
@@
Figure 11: Failure mode distribution across eight experimental conditions (n=20n{=}20 each) in the pharmaceutical domain. Baselines exhibit diverse failure modes dominated by impurity and conversion violations. Mono Opus thinking (Cond. 5) exhibits a dual-attractor pattern (C1+C2 optimum vs. C4 boundary). Mono+UAI (Haiku true tool-call, Cond. 7) enters the stochastic-oscillation trap with all trials hitting max_iters (Finding 8). CAAF achieves perfect paradox detection in both Cond. 3 (all-mini) and Cond. 8 (all-Haiku).
@@ 7.3 Analysis
Finding 7 — Hint Fragility Scales with Complexity. In the AD domain (2-constraint paradox), the hint condition achieved 100% detection. In the pharma domain (3-way paradox with nonlinear interactions), hint-aided detection drops to 70%. This confirms that prompt engineering reliability degrades as constraint complexity increases—precisely the regime where architectural enforcement becomes most critical.
Finding 8 — Three-Tier UAI Ablation: UAI Access is Necessary but Not Sufficient. The Mono+UAI condition strips away RAD and Structured Semantic Gradients, giving a monolithic LLM direct UAI access. We evaluate three implementations at n=20n{=}20 on Pharma: (i) prompt-simulated UAI on GPT-4o-mini (Cond. 4), where the model is told to mentally apply the harness formulas and retry up to three times; (ii) true tool-call UAI on Claude Haiku 4.5 without extended thinking (Cond. 7), where the model invokes a Python harness assertion engine via the Anthropic tool-use API at commodity-tier reasoning cost; and (iii) true tool-call UAI on Claude Opus 4 with adaptive high-effort thinking (Cond. 6), same tool interface at frontier-tier reasoning cost.
Prompt-simulated UAI on the commodity model reaches 95%; true tool-call UAI on the frontier model reaches 100% (all 20 trials correctly identifying the minimal unsatisfiable subset $\{C1,C2,C4\}$ in a mean of 3.15 tool-loop iterations); but true tool-call UAI on the commodity model reaches 0%. In Cond. 7, 20/20 trials exhaust max_iters=8 oscillating between CONVERSION_MINIMUM and IMPURITY_LIMIT (final-iteration failed-rule tally: 14 IMPURITY, 12 CONVERSION, overlapping), and no trial declares a paradox. The variance across the three cells is therefore not explained by UAI access modality: the configuration with the most faithful UAI access (case ii—every constraint evaluated by a Python engine, every verdict returned as a structured tool_result) performs worst. It is explained by the interaction between reasoning capability and the capacity to step back from iterative local fixes to recognize global unsatisfiability. Case ii enters the stochastic-oscillation trap identified in §6 for naive-reflection baselines: it survives transplantation into a Mono+UAI scaffold because Mono+UAI provides detection but no State Locking—each tool-call turn repairs one constraint while re-introducing another.
[…]
@@
  • UAI (Harness as an Asset) provides the deterministic constraint boundary—necessary for grounding but not sufficient alone at commodity reasoning cost.
@@
  • Structured Semantic Gradients with State Locking prevent the stochastic-oscillation trap that destroys Mono+UAI with commodity models; this pillar is reliability-load-bearing at the price point where deployment economics become practical.
[…]